How do I approach this?61. Use Eulers method to find approximate values for the solution of the initial value problem dy dx = I – Y y(0) 1 on the i
Answer #1

y=t-y 910)=1. h=0.2=1/50 (replace Step 1. to=otlls = 115. (2=t)}. y (tu) = y(15) - y, - yothf(to yo) = 1th Ho, 1) =0.8 Il fa